From: STOSC::SMM "On-Duty SMM x5074" 20-JAN-1993 07:53:45.53 To: @[.DIS]HST_STAT CC: Subj: HST Status Report, 01/15 - 01/19. HUBBLE SPACE TELESCOPE DAILY REPORT #805 PERIOD COVERED: 0800 EST 01/15/93 - 0800 EST 01/19/93 Daily Status Report as of 019/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera was used to take a series of high latitude images which were taken in parallel with the prime science observations. All of the observations were executed as scheduled, and the observations were reported to to be full of stars. 1.2 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.3 Completed GHRS/2 3372 (Point Spread Function Determination) The High resolution Spectrograph was used to make a large series of observations of the point spread function (PSF) as it relates to the large science aperture (LSA). Observations were made using the standard stellar target HD141556. All the observations executed with no apparent problems in the telemetry. 1.4 Completed PC 4140 (Rapid Internal Monitor - Part 2) The Planetary Camera was used to take internal flat observations to check for new forms of contamination and the return of quantum efficiency hysteresis (QEH). The observations were monitored in the OSS, and no problems were reported. 1.5 Completed FOS/BL 4057 (Spectrography of QSOs) The Faint Object Spectrograph (Blue Side) was used to make three ultraviolet spectral observations below 3000 A0 of quasar I-ZW1. The execution of the observations were monitored in the OSS, and all of the observations were reported to look okay. 1.6 Completed FOS/BL 4110 (The Physics of Massive O-Type Stars) The Faint Object Spectrograph was used to make set of nine ultraviolet (from 1150 A0 to 2300 A0) spectral observations of OB class star SK67D166 which is located in the Large Magellanic Cloud. The first six observations were received and displayed. They all looked good. The next onboard binary target acquisition succeeded, but the following peak up had the maximum count in the last group, which indicates that the centering of the target was not well known. The last observations were reported to look good. 1.7 Completed WFC-PC 4109 (Rapid Internal Monitor) Both the Planetary Camera and the Wide Field Camera using filter F555W were used to take two internal flat observations to check for new forms of contamination (i.e., Daisies and measles). Both images were received and displayed in the OSS. They both were reported to look normal. 1.8 Completed Two Sets of PC 4714 (Rapid Internal Monitor) The Planetary Camera was used to make two internal flat observations. The images will be used for the restoration of Earth flats, which have been altered due to changes in the quantum efficiency which occurred after WFPC decontamination operations. Both observations were executed without apparent problems, and were reported to look normal. 1.9 Completed FOS/RD/BL 3647 (Star forming Histories of Elliptical Galaxies) The Faint Object Spectrograph was used to make two spectral observations (both the blue and red sides) of elliptical galaxy NGC-1399. This is to study the constraints on the metallicity and age distribution of the hot star in this galaxy. Part of the first observation was not received due to a TDRSS hand-over which interrupted the data transmission. The other observations looked noisy, but had enough signal after the observations had been co-added. 1.10 Completed WFC 4105 (HST Medium Deep Survey High Latitude) The Wide Field Camera was used to make two observations to be used for a medium-deep sky survey. These observations were take in parallel with the primary science observations. There were no apparent problems. 1.11 Completed Two Sets of WFC 3917 (HST Medium Deep Survey - Part 1) The Wide Field Camera was used to make two observations which were taken in parallel with the prime science observations. The observations were monitored in the OSS, and no problems were reported. 1.12 Completed FOC/96 3880 (Absorption Lines in Planetary Nebulae) Observations were made of the central star in protoplanetary nebula BD+30D3639. Star light from the central star is used to probe the neutral material surrounding this star. Two observations were executed in fine lock. A recentering event occurred during the first observation and had no effect on the observation. Both images were reported to have stellar type object within the field of view. 1.13 Completed WFC 1306 (GTO Low Latitude Parallel Program) The Wide Field Camera was used to make two low latitude observations taken in parallel with the prime science observations. These observations are looking for low mass stars, close visual binaries, and shells surrounding late type stars. Both observations were executed in fine lock. There was no loss of lock, but a recentering event did occur. Both images were reported to be full of stars. 1.14 Completed FOS 4000 (Quasar Absorption Line Survey) The Faint Object Spectrograph was used to make series of observations for a spectral line survey of Quasar PKS2340-03. There were binary and peakup acquisitions followed by a RAPID-mode spectra using the FOS red side detectors. The observations were executed as planned, and they encountered no apparent problems. 1.15 Completed FOC 3801 (Search for QSOs Suitable for Subsequent Observations) The Faint Object Camera was used to make ultraviolet flux observations of quasar 0118+0119 to search for quasars to be used in further detailed studies. The observations were not monitored in real time. After the science tape recorder playback, the two images showed no visible objects. Since this was one of the possible expected outcomes of the observations, the observations were not reported as being bad. 1.16 Completed FOC 3815 (UV Observations of Hubble Sandage Variables) The Faint Object Camera was used to make two images of Hubble-Sandage class stars which are located in galaxy M-31 (aka, the Andromeda Galaxy). These are very luminous, eruptively unstable stars of the same type as Eta Carina. The science tape recorder playback of the image showed multiple targets in both images 1.17 Completed PC 4764 (Investigation of Dynamics of Binary Asteroids) The Planetary Camera was used to take a series of images of suspected binary asteroid 18 Melpome. Its binary nature has been suspected because of its unusual light curve. The baseline guide star acquisition was made in course track. The profiled slew and the image observations of the asteroid were successful. 1.18 Completed Three Sets of FOS/RD 3744 (Spectrophotometry of Phobos and Deimos) The Faint Object Spectrograph (Red Side) was used in the prism mode with the one arcsecond aperture to make observations of Mars' two moons Phobos and Deimos. The initial guide star acquisition failed, and as preplanned the FOS was safed by stored command load instructions at 018/1054Z. The FOS was recovered by real-time commanding at 018/1815Z. The recovery was successful. The following baseline acquisition for observations of the moon Phobos were successful. There was some trouble holding the image in the observation aperture window. A corrective SIP maneuver was uplinked. The onboard target acquisition succeeded and executed with no problems. During a subsequent observation, the baseline guide star acquisition for observations of Deimos failed. The FGS #2 had a stop flag and search radius limit exceeded flag set. A corrective SIP maneuver was uplinked, however, following observations did not show the moon Deimos in the field of view. 1.19 Completed PC 3603 (Snapshots of Protoplanetary Nebulae) The Planetary Camera was used to make a single observation of the protoplanetary nebula around this star PK211-03D1. The observations was received and displayed. It was reported to look good. 1.20 Completed FOS/RD/BL 4694 (Stellar Content of Galaxies and Globular Clusters) An attempt was made to make observations of four different globular clusters in the Galaxy M-31 (aka, Andromeda Galaxy). However, The planned six observations did not occur because the FOS was safed. An attempt on the last observation was bad because the onboard target acquisition did not occur. 1.21 Completed PC 4106 (HST Medium Deep Survey: Low Latitude) The Planetary Camera was used to make a series of observations taken in parallel with the prime science observations. These are to provide for serendipitous discoveries and to supply data for a basic sky survey. The observations were executed as scheduled, and no problems were reported. 1.22 Completed PC 4712 (Measles Monitor) The Planetary Camera was used with filter F517N to obtain Earth flat calibration images to monitor the appearance and number of measle contamination points. The observations were monitored in the OSS, and no problems were reported. 1.23 Completed PC 3230 (PSF and Focus Monitoring 2) The Planetary Camera was used to observe the 8th magnitude standard star SAO55685 to monitor the point spread function and focus changes caused by desorption and other changes within the WFPC system. The observations were executed in course track. All the data was received and displayed. In one of the images the star appeared on the Baum spot, and the observation was reported to be bad. 1.24 Completed HSP/UV2 3798 (Search for Radial Pulsations in White Dwarfs) An attempt was made to look for radial pulsations in the hot white dwarf star G130-5 using the High Speed Photometer. The baseline guide star acquisition failed to obtain fine lock. The stop flag and the search limit exceeded flag were set for FGS #2. The observations was monitored in real-time, and there were almost no counts in the FGS #2 PMTs. 1.25 Completed FOS/BL 3816 (White Dwarf Stars) The Faint Object Spectrograph was used to obtain ultraviolet spectral observations of white dwarf star G29-38. The observations were monitored in the OSS, and no problems were reported. 1.26 Initiated HSP/UV2 1095 (Variability of High Luminosity Stars) Observations have been initiated using the High Speed Photometer to monitor one of the most massive and luminous known stars HD15570. The observations are currently under way, and no problems have been reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 19 Successful : 16 (CT at 018/1037Z, and 2328Z, FL at 019/0216Z) Scheduled Reacquisitions : 41 Successful : 40 (FL at 019/0352Z) Losses of Lock : 0 2.2 FHST Updates: Scheduled : 36 Successful : 36 2.3 Significant HSTARs: HSTAR 3810 The Faint Object Spectrograph was automatically placed in Safe Mode by a stored command load at 018/1054. This was during an observations of Mars' two moons. The baseline guide star acquisition for observation 3744 failed, and as planned the FOS was placed in Safe Mode to prevent an over-light condition in the FOS. The FOS was recovered by real-time commanding at 018/1815Z and subsequent observations using the FOS were successful. (See Section 1.18) . 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete HSP/UV2 1095 (Variability of High Luminosity Stars) 3.2 Complete GHRS/2 3737 (Onset of Chromospheric Activity) 3.3 Complete WFC-PC 4109 (Rapid Internal Monitor) 3.4 Complete PC 3230 (PSF and Focus Monitoring 2) 3.5 Complete Two Sets of GHRS/2 3578 (Line Eclipse Mapping of Accretion Disk Winds) 3.6 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.7 Complete One Set of FOS/RD/BL 4694 (Stellar Content of Galaxies and Globular Clusters) and Initiate a Second 3.8 Complete FOS/BL-WFC 3545 (Evolution of UV Spectra in Early Type Galaxies) 3.9 Complete FOS/RD 4052 (Inner Regions of Quasars) 3.10 Complete One Set of WFC 3917 HST Medium Deep Survey: Part 1) and Initiate a Second 3.11 Complete WFC 4105 (HST Medium Deep Survey: High Latitude) 3.12 Complete Circumstellar and Interstellar Absorption Lines) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-